A smoke plume is rising through an atrium at the rate of a metre per second. Calculate its momentum, if its density is 0.77 kg/m3.
Answer #1

Momentum (M) of smoke plume will be given as:
M=m*u
where m=mass flow rate of smoke plume=density×Volume flow rate=0.77×1=0.77 kg/s per m3/s (assuming unit volume flow rate)
u=Velocity of rise=1 m/s
Therefore,
M=0.77×1=0.77 kg*m/s per m3/s